OpenStudy (anonymous):

Create your own binomial expression with a radical in the second term. Part 1: Identify the conjugate and explain, in complete sentances, why it is the conjugate. (1 point) Part 2: Multiply your original binomial expression and its conjugate. What happened to the radicals and why? (1 point) Last one for right now, peopleee <3 please!

OpenStudy (anonymous):

So now, how do i solve it in reguards to my directions?

OpenStudy (anonymous):

To identify the conjugate acid, look for the pair of compounds that are related

OpenStudy (mathstudent55):



OpenStudy (anonymous):

I really just need the answer in step form. like step 1: asdfjak;dlfj Step 2: ajkfak;dfg

OpenStudy (anonymous):

1+√5

OpenStudy (anonymous):

This is with the sign changed though 1 - √5

OpenStudy (anonymous):

(1-√5)(1+√5) = 6

OpenStudy (mathstudent55):

The conjugate has the same rational part (the 5 in this example) and the opposite sign on the radical part (sqrt(3) in this example). The conjugate is 5 - sqrt(3)

OpenStudy (anonymous):

2. The radicals disappear because you're using the difference of two squares: (a-b)(a+b) = a^2-b^2 and a radical squared is rational

OpenStudy (anonymous):

omg thank you <3